OpenStudy (anonymous):

The height H in feet above the ground of the seat of a playground swing can be modeled by H(θ) = -4cosθ + 6, where θ is the angle the swing makes with a vertical extended to the ground. Graph the height of the swing's seat for 0º ≤ θ ≤ 90º. Then answer this question: How high is the swing when θ = 60º?

OpenStudy (anonymous):

do we draw a acute triangle for this problem?

OpenStudy (anonymous):

do you have a graphing calculator?

OpenStudy (anonymous):

Simply substitute 60 degrees into the equation H(60) = -4cos(60) +6 = -4*0.5 + 6 =-2+6 = 4 units up.

OpenStudy (anonymous):

do what he said but as for the graphing portion just plug it into a graphing calculator and make the x-min: 0 and x-max: pi/2 since 90 degrees is pi/2